Use trigonometric identities to solve the equation 2sin(2θ)-2cos(θ)=0 exactly for 0≤θ≤2π.

A.) What is 2sin(2θ) in terms of sin(θ)and cos(θ)?

B.) After making the substitution from part 1, what is the common factor for the left side of the expression 2sin(2θ)-2cos(θ)=0 ?

C.) Choose the correctly factored expression from below.

a.) cos(\Theta )(4sin(\Theta )-2)=0

b.) 4 sin (\Theta ) cos(\Theta ) = 2 cos (\Theta )

c.) 2 cos (\Theta ) 2 sin (\Theta ) = 0

d.) 2cos(\Theta )(2sin(\Theta )-1) = 0
Answer #1

Answer: Given equation 2 Sin (20)-20050 =0 a) given that 2 sin (20) in terms of sin(o) and cosco) According to formula sin(2A2 Sin (20)-2 COJCO)=0 4 sinco) Cos (0)-2 Cos (0) 50 2 cos (o) [2 Sin (0)-1] =0 common factor for the left side of the express